The following data represents the students' scores :


(81,35,86,91,69,72,47,66,80,55,90,53,43,56,80,77,59)


Wanted to find :


1_SMA 2 Mediator 3_ Vein


4_Term 5_standard deviation 6_Variance


1
Expert's answer
2021-06-08T07:28:55-0400

Given data

81,35,86,91,69,72,47,66,80,55,90,53,43,56,80,77,59



(1)SMA (Simple Moving Average)


data: 81,35,86,91,69,72,47,66,80,55,90,53,43,56,80,77,59

order wise:

Number of Consecutive Points to Average: 2 {for first SMA , it is fixed is

2}

SMA="\frac{81+35}{2},\frac{35+86}{2},\frac{86+91}{2},...\frac{80+77}{2},\frac{77+59}{2}"

hence we get

SMA=58,60.5,88.5,80,70.5,59.5,56.5,73,67.5,72.5,71.5,48,49.5,68,78.5,68


______________________

(2)mediator

rearrange data

such date

35, 43, 47, 53, 55, 56, 59, 66, 69, 72, 77, 80, 80, 81, 86, 90, 91

hence mediator value is 69.

__________________

(3) it is meaning of the word is missing in statistics and probability. so we do not provide a solution of it .

______________________

(4)given data

81,35,86,91,69,72,47,66,80,55,90,53,43,56,80,77,59

then the number of terms is equal to 17.

____________________

(5)_standard deviation

"\sigma=\sqrt{\frac{1}{N} \sum_{i=1}^{N}\left(x_{i}-\mu\right)^2}"

"\sigma =\sqrt{\frac{(81 - 67.058823529412)^2\\ + ... + (59 - 67.058823529412)^2
}{17}
}\\
=\sqrt{\frac{4654.9411764706}{17}}\\
=16.547509456234\\"

______________________

(6)Variance


"variance =\sigma^2\\\sigma^2 =(16.547509456234)^2\\
=273.82006920415"